OpenStudy (anonymous):

Solve 7^x – 5 = 7^4x + 7

mathslover (mathslover):

is it : \[\large{7^{x-5}=7^{4x+7}}\]

OpenStudy (anonymous):

Yes it is

mathslover (mathslover):

ok so since bases are same powers will be equal that is x-5 = 4x+7

mathslover (mathslover):

can u solve for x now?

OpenStudy (anonymous):

so x=4x+13 now i divide 4 correct?

mathslover (mathslover):

mistake there... check it up 7+5 is not equal to 13

OpenStudy (anonymous):

im sorry I meant x=4x+12 so now I divide by 4 correct?

mathslover (mathslover):

no don't divide by 4 subtract 4x both sides

OpenStudy (anonymous):

okay so I got -4x= 12 now what should i do?

OpenStudy (anonymous):

I believe I divide -4?

mathslover (mathslover):

I think you did mistake again : \[\large{x=4x+12}\] \[\large{\color{violet}{x-4x=4x-4x+12}}\] \[\large{\color{red}{-3x=12}}\]

mathslover (mathslover):

nod divide by -3

OpenStudy (anonymous):

so would my answer be x=-4?

mathslover (mathslover):

yes right

OpenStudy (anonymous):

Thank you so much!

mathslover (mathslover):

welcome dear also: \[\huge{\cal{Welcome}\space \mathsf{To}\space \cal{ openstudy}}\]